# Unstable Unlearning: The Hidden Risk of Concept Resurgence in Diffusion Models

### v1

@@ 3 Composing Updates Causes Concept Resurgence
Following [31], the second benchmark we consider is inducing the model to forget how to generate certain types of objects from the CIFAR10 dataset (the “erase set”) while retaining the ability to generate others (the “retain set”). We apply MACE to Stable Diffusion v1.4 across four subtasks. In the first three, we apply MACE to erase knowledge of a single object (automobiles, ships and bird, respectively). In the fourth, we apply MACE to a more challenging task in which the goal is to simultaneously unlearn a set of five concepts (automobiles, ships, birds, cats, and trucks). In each subtask, we then evaluate whether the model can generate images of these objects and their synonyms (e.g., after being prompted with “a photo of the [erased object]”). Both the full set of erased objects and retained objects, along with the prompts used to evaluate the model, are provided in Section C.1. As in the celebrity erasure task, we adopt the set of concepts to be erased, evaluation prompts and other hyperparameters from [31].222The only exception is the Erase 5 Objects task, which we add to evaluate simultaneous erasure of multiple concepts. We quantify model performance by computing the CLIP accuracy across three random seeds on the set of evaluation prompts. Following [31], we do not compute FID-5K and CLIP-5K for the object erasure task; COCO-5K is itself composed of common objects, and the goal of this evaluation is to assess performance on generic concepts unrelated to the erasure task.
As discussed above, object erasure and celebrity erasure are two of the four tasks considered in [31]. We discuss our choice to exclude the other two (artistic style and explicit content erasure) in Appendix A.
Evaluating concept resurgence.
We present representative examples to characterize the degree of “typical” concept resurgence in Figure 2 and Figure 3, and curate specific examples of this vulnerability in Figure 4.333We say these are “representative” examples because we choose the first image generated for each prompt after fixing the random seed at , and use the same set of unlearned concepts, hyperparameters and prompts studied in [31]. Thus, these figures highlight the degree of concept resurgence in a “typical” case.
As Figure 2 demonstrates, concept resurgence can occur in degrees, as some concepts are not reintroduced at all (e.g., Melania Trump), and others are only partially reintroduced (e.g., Barack Obama). Furthermore, Figure 3 demonstrates that concept resurgence can be rare in some contexts; indeed, none of the representative images we sample reintroduce the unlearned concepts in the object erasure task. However, as both Figure 2 and Figure 4 demonstrate, concept resurgence can occur in striking and seemingly unpredictable ways, running the risk that developers or users can inadvertently reintroduce harmful or unwanted content.
In Figure 5, we quantify the degree of resurgence across both the object and celebrity erasure tasks using the metrics described above. As suggested by the qualitative results, the degree of resurgence is substantially larger in the celebrity erasure task, particularly as the number of unlearned celebrities grows large (intuitively, such tasks are “harder” than unlearning a smaller number of celebrities, as the model must simultaneously unlearn many concepts without degrading model performance on other, unrelated tasks). The vulnerability appears in the object erasure task as well, albeit to a lesser degree. We emphasize however that in many contexts, even rare concept resurgence presents unacceptable risks. In the remainder of this work, we seek to systematically characterize when and why this resurgence occurs, focusing first on algorithmic choices (Section 4) and then on characteristics of the fine-tuning dataset (Section 5).
4 Algorithmic Factors Driving Concept Resurgence
The compositional updating pipeline involves several algorithmic choices that can either contribute to or mitigate the risk of concept resurgence. We focus on common choices made across most concept unlearning algorithms for diffusion models, using MACE as our baseline technique to assess their impact. These choices include regularization in the cross-attention refinement update, mapping the embedding of the concept to be unlearned to more general concepts, and the selection of the algorithm used for fine-tuning. In the following two sections we focus our attention on the erase 10 celebrities and erase ship tasks as we vary other components of the training pipeline.
At a high level, we hypothesize that resurgence occurs because unlearning does not update the model parameters to be sufficiently “far away” from the pretrained weights. Thus, although unlearning may initially suppress the generation of unwanted concepts, even a modest degree of fine-tuning tends to shift model weights towards their initial state, and thus reintroduce seemingly erased concepts. To validate this hypothesis, we first assess the impact of three algorithmic factors on concept resurgence: choice of mapping concept, regularization, and the fine-tuning algorithm.
4.1 Mapping Concepts
A key algorithmic component in many approaches to concept unlearning in diffusion models, including MACE and Unified Concept Editing (UCE) [13] is modifying the cross-attention mechanism. This mechanism is responsible for encoding the prompt into an embedding that the diffusion process conditions on. In particular, MACE proposes to modify the cross-attention weights found via the solution to the following optimization problem:
where is the set of pretrained weights, is the embedding of the word that co-exists in the prompt with the concept to be unlearned (e.g. the photo of {concept}), and is the embedding of the word that co-exists with the concept to be unlearned if that concept was replaced with its more generic concept (e.g. the photo of {generic concept}). Finally, is an embedding for a concept that we would like to preserve. Typically, the preserved concepts are generic ones that the model should still be able to generate after unlearning. The second term acts as a regularizer that keeps the new weights close to the original weights and thus prevent performance degradation on unrelated concepts.
In this section we focus on the choice of the generic mapping concept. This impacts both the cross-attention refinement objective and the final update applied to the model, which fuses multiple LoRA modules corresponding to each unlearned concept. As described in Eq. 1, the embeddings of the words which co-occur with the concept to be unlearned are mapped to the corresponding embeddings for the mapping concept. To test our initial hypothesis, we select mapping concepts which are increasingly unrelated to the concept to be unlearned. This acts as a proxy for moving the initial model farther and farther away from its initial pretrained weights. For erase 10 celebrities we experiment with the following mapping concepts: “a person” (the baseline), “an object” and “a male, female celebrity,” the latter two of which correspond to more general and more specific concepts, respectively. For erase ship we experiment with “ground” (the baseline) and “object.” We present these results in Figure 6.
As illustrated in Figure 6, concept resurgence is quite sensitive to the choice of mapping concept. In the object erasure task, mapping “ship” to a different but specific concept (“ground”) reduces resurgence more effectively than mapping it to a general concept like (“object”). In the celebrity erasure task, mapping each celebrity to a more specific concept (e.g., Jennifer Aniston “a female celebrity”) leads to more severe resurgence than if erased celebrities are mapped to more generic concepts. These findings are consistent with the hypothesis that while unlearning can initially suppress the generation of the unlearned concepts, it may do so through small changes in the parameter space, which can be easily undone by even modest degrees of further fine-tuning.
4.2 Regularization
In this section we study the impact of the regularization parameter (recall Equation 1) and assess its impact in the degree of concept resurgence. We run the MACE algorithm with in the erase 10 celebrities and erase ship tasks. Intuitively, small values of allow for large model updates, and thus prioritize unlearning the unwanted concepts at the expense of possibly degrading overall model performance. In contrast, large values of only allow the unlearning algorithm to make small perturbations to the initial model weights.
As Figure 7 demonstrates, the degree of regularization is correlated with the degree of concept resurgence, particularly in the celebrity erasure task. Once the regularization parameter surpasses we see that more than half of the images generated in the forget set are accurately classified as their concept, indicating that larger updates to the initial pretrained weights are needed to ensure the prevention of concept resurgence. In the object task, we find that the regularization does not impact concept resurgence. Based on our results in Sec. 4.1, it is clear that the impact of mapping concepts on the phenomena of concept resurgence is first order. The object erasure task uses the mapping concept ground which is completely unrelated to ship. Celebrity erasure uses person as the mapping concept, which is much more related. Decreasing regularization helps prevent concept resurgence when the mapping concepts chosen are more related to the unlearned concepts.
4.3 Fine-Tuning Algorithm
Finally, we investigate the impact of fine-tuning the unlearned models with low rank adaptation (LoRA) instead of traditional full-parameter fine-tuning algorithms. LoRA is the most common method for fine-tuning large scale diffusion models (like Stable Diffusion v1.4). Although full parameter fine-tuning is far less efficient than LoRA, and thus not always feasible to implement in practice, this experiment seeks to investigate the effect of full parameter fine-tuning on concept resurgence. In particular, full parameter fine-tuning will tend to make larger incremental updates to the model parameters than LoRA under otherwise similar starting conditions.
Fig. 8 demonstrates that concept resurgence is largely insensitive to the choice of fine-tuning algorithm, and, if anything, that full fine-tuning can exacerbate the degree of concept resurgence in some cases. This provides additional evidence that vulnerability to concept resurgence is largely due to the initial unlearning algorithm rather than choices made at the fine-tuning stage. In particular, these results, in conjunction with the regularization and mapping concept experiments above, lend support to the hypothesis that larger parameter updates are required in the unlearning phase to prevent concept resurgence.
5 Data-Dependent Factors Driving Concept Resurgence
The second component of the pipeline we investigate is the data the unlearned models are fine-tuned on. As discussed in Section 3, we choose not to focus on adversarial dataset constructions (including e.g., simply directly fine-tuning a model on concepts which were previously unlearned); We refer the reader to [33, 44, 9] for works on attacking or jailbreaking text-to-image diffusion models. Instead, we consider the kind of dataset constructions that occur as part of common and benign use. For example, the end user of an open source diffusion model may want to fine-tune the model to acquire new concepts which were excluded from the pretraining set and/or improve its performance on particular tasks of interest. In Section 3, we demonstrated that fine-tuning on random in-domain concepts (e.g., fine-tuning on randomly chosen celebrities that were not part of the erased set) can lead to concept resurgence. In this section, we seek to further investigate the role of dataset construction in concept resurgence. First, we consider fine-tuning on in-domain concepts with varying levels of similarity to the concepts which are initially unlearned. Intuitively, it may be that fine-tuning on more related concepts can exacerbate the degree of concept resurgence. We then turn to fine-tuning on out-of-domain concepts; e.g., fine-tuning on images of randomly chosen objects after a celebrity unlearning task. We describe these experiments in more detail below.
5.1 CLIP Distance
First, we consider fine-tuning on unrelated but in-domain concepts as described above. We further segment these concepts by thresholding the CLIP distance to the unlearned concepts, which we use as a proxy for how “related” the fine-tuning dataset is to the unlearned concepts. In particular, for each task we find a publicly available dataset with hundreds of concepts in the same domain (e.g., for the celebrity unlearning task, this is a dataset of celebrity images). We describe these datasets in detail in Section C.2. For each concept in the corresponding dataset, we compute its minimum CLIP cosine similarity over the set of unlearned concepts. We then partition this fine-tuning set into three evenly sized subsets based on the percentiles of the CLIP cosine similarity. Finally, we randomly sample 10 concepts from each tercile to create different fine-tuning datasets which vary in their degree of “relatedness” to the unlearned concepts. We present these results in Figure 9.
As Figure 9 demonstrates, the degree to which the fine-tuning set is related to the unlearned concepts — at least as measured by CLIP cosine similarity — does not appear to meaningfully correlate with the degree of concept resurgence. This finding underscores the danger of concept resurgence even when fine-tuning on relatively unrelated data.
5.2 Out-of-Domain Concepts
Finally, to better understand the scope of concept resurgence, we curate an additional set of fine-tuning datasets which contain “out-of-domain” concepts which are wholly unrelated to those in the unlearning task. In particular, for each unlearning task, we fine-tune the resulting model on the random concept datasets from the other two unlearning tasks (e.g., for the celebrity unlearning task we further fine-tune the resulting model on the same randomly selected objects used as the initial fine-tuning set in the object erasure task).
As Figure 10 demonstrates, the degree of concept resurgence does vary with the domain of the fine-tuning set, particularly for the celebrity task. In particular, fine-tuning on in-domain images appears to exacerbate the risk of concept resurgence, while fine-tuning on out-of-domain images can mitigate it. Thus, although we found in the previous section that the degree of concept resurgence is relatively insensitive to variation in the fine-tuning set within each domain, it appears that fine-tuning on images which are wholly unrelated to the unlearned concepts can be safer than fine-tuning on more closely related images. This finding thus lends additional support to the hypothesis laid out in Section 4.
@@ 6 Discussion and Limitations
The scale of generative models introduces new challenges, including the risk of learning concepts that are unsuitable or undesirable for certain downstream applications. Ideally, unlearning methods would allow model developers to precisely and permanently remove unwanted concepts while preserving the model’s overall utility. Reality, however, is more complex.
@@
Our investigation opens up several important avenues for future work. For example, we do not provide a theoretical characterization of concept resurgence, nor do we present a strategy designed to prevent it from happening. Both developments could help to enhance the robustness of unlearning methods. Additionally, though our evaluations focus on well-known academic benchmarks, further research is necessary to assess the prevalence of concept resurgence in real-world deployments (particularly the effect of interleaving a large number of compositional updates, which may exacerbate the vulnerabilities we demonstrate here).
@@
Acknowledgements
A.W. was funded in part by the Lister Brothers Career Development Professorship at MIT. V.S was funded in part by an MIT GenAI impact grant. A.S. acknowledges support from the Simons Foundation and NSF through award DMS-2031883, as well as from the DOE through award DE- SC0022199. R.A. and M.R acknowledge support from A Stephen A Schwarzman College of Computing Seed Grant.
Appendix A Excluding Artistic Style and Explicit Content Erasure
As discussed in Section 3, celebrity and object erasure are two of the four benchmarks considered in [31]. The other two considered are artistic style erasure — unlearning the ability to generate images in the style of specific artists, e.g., due to copyright concerns — and explicit content erasure, particularly to suppress nudity. We exclude artistic style erasure due to the difficulty of quantifying the degree of a particular style in an image. In particular, unlike celebrity or object erasure, artistic style is not localized to specific regions in an image, and instead is a holistic (and partially subjective) property of the model output. For example, [10] note that “courts have emphasized the importance of considering the ‘total concept and overall feel’ [for determining whether copyright infringement has occurred], rather than relying on mechanical dissection or quantitative measures alone.”
Characterizing artistic style replication (and copyright infringement more broadly) is rich topic in its own right, and we refer to [42, 40, 46, 6] for additional background.
We further exclude the explicit content benchmark due to the sensitive and unpredictable nature of the images which may be generated by the model, the lack of agreed upon standards for conducting such evaluations responsibly, and recent well publicized examples of the real-world harm that can result from synthetic but realistic nude images [26, 15]. Instead, we use the object and celebrity erasure tasks as representative but benign benchmarks on which to conduct our evaluations.
Appendix B Impact of Fine-tuning on Retained Concepts
Below we first examine the analogue of Figure 5 on the retained set, which is presented in Figure 11. Consistent with [31], we find that MACE preserves model performance on the set of retained concepts, and furthermore, subsequent fine-tuning does not degrade performance on this set.
As an additional sense check, we examine the CLIP and FID scores on random objects sampled from COCO-5K (as described in Section 3) before and after fine-tuning in the celebrity erasure tasks. These are presented in Figure 12 and Figure 13, respectively. We see that the CLIP scores remain almost identical, while the FID scores increase (i.e., degrade) after fine-tuning. The results of these three figures are thus broadly consistent with fine-tuning not degrading performance across a variety of tasks; if anything, concept resurgence can occur even if overall performance (i.e., on unrelated tasks) decreases slightly.

### v2

@@ 3 Composing Updates Causes Concept Resurgence
Object erasure. Finally, following [27], the final benchmark we consider is inducing the model to forget how to generate certain types of objects from the CIFAR10 dataset (the “erase set") while retaining the ability to generate others (the “retain set"). We apply each unlearning algorithm to Stable Diffusion v1.4 to erase three objects (automobiles, ships, and birds) simultaneously. We then evaluate whether the model can generate images of these objects and their synonyms (e.g., after being prompted with “a photo of the [erased object]"). Both the full set of erased objects and retained objects, along with the prompts used to evaluate the model, are provided in Appendix C. As in the celebrity erasure task, we adopt the set of concepts to be erased, evaluation prompts and other hyperparameters from [27].222The only exception is the Erase 5 Objects task, which we add to evaluate simultaneous erasure of multiple concepts. We quantify model performance by computing the CLIP accuracy across three random seeds on the set of evaluation prompts.
Evaluating concept resurgence. In each of these settings, we are primarily concerned with whether concept resurgence occurs, and, if it does, the rate at which it does so. We curate specific examples to characterize the severity of concept resurgence in Figure 2. We show concept resurgence can occur in striking and seemingly unpredictable ways across all four algorithms, running the risk that developers or users can inadvertently reintroduce harmful or unwanted content.
In Table 1 and 2, we quantify the degree of resurgence across all four tasks and unlearning algorithms using the metrics described above. The degree of resurgence varies across the algorithms and tasks. ESD and SDD exhibit a large degree of concept resurgence across all tasks; in some cases benign fine-tuning reverses unlearning almost completely. For MACE we see a modest degree of concept resurgence across all four tasks, and for UCE we see a small amount of resurgence in the celebrity and object erasure tasks. These findings illustrate that concept resurgence occurs with striking regularity across both algorithms and domains. We emphasize that in many contexts, even rare concept resurgence presents unacceptable risks. In the remainder of this work, we characterize the factors that affect the severity of concept resurgence and investigate the root causes of this phenomenon.
4 Factors Influencing Concept Resurgence Severity
We find two important components of the compositional updating pipeline that influence the severity of concept resurgence. The first is the number of concepts that were simultaneously unlearned. The second is the techniques used in the unlearning algorithms.
4.1 Scaling Unlearning Algorithms
A key desideratum for any unlearning algorithm is the ability to scale: ideally, the user can erase many concepts without retraining the model from scratch. All four unlearning algorithms we consider report the ability to simultaneously unlearn many concepts while maintaining utility on unrelated concepts. We analyze whether increasing the number of concepts which are unlearned leaves the resulting model more susceptible to concept resurgence. For the celebrity erasure task, we define four subtasks: erasing , , , and celebrities. For the object erasure task, we define three subtasks: erase ship, erase three objects (automobile, ship, bird), and erase five objects (automobile, ship, bird, cat, and truck). We follow the same evaluation setup as described in Section 3 for both tasks. We omit the copyright task from this analysis because we found that the models were unable to unlearn more than one character without dramatically degrading performance on retained characters.333In this case, we interpret the algorithm as having failed in the first unlearning step, and thus there is no potential resurgence to evaluate. Without this requirement, a model which simply outputs random noise would suffice to achieve perfect performance on any unlearning task. We also omit the unsafe content task, as it cannot be cleanly decomposed into discrete “subtasks" (e.g., individual celebrities, objects or characters).
The impact of increasing the number of unlearned concepts varies amongst the four algorithms. For ESD, there is clear increase in resurgence as the number of concepts unlearned increases (Figure 3). In contrast, for MACE, UCE, and SDD the level of resurgence was not impacted as the number of concepts increased (see Appendix E). We discuss the possible mechanisms at play in the following section.
4.2 The Impact of Algorithmic Choices on Resurgence
The four algorithms we consider perform unlearning through fine-tuning model parameters, closed-form edits, or a combination of both. Fine-tuning optimizes an unlearning objective via gradient-based methods, as seen in ESD, which adjusts the model so that the score function conditioned on a concept matches the unconditional score function. Closed-form edits derive an explicit update for unlearning, as in UCE, which modifies key and value weights in cross-attention layers to replace concept-specific representations with generic or blank ones. MACE combines both approaches: it uses a closed-form edit to adjust word embeddings in concept-containing prompts and LoRA fine-tuning to suppress concept-related attention in generated images. We categorize ESD and SDD as fine-tuning methods, UCE as closed-form, and MACE as a hybrid approach.
Finetuning vs. Closed-Form In Table 1, we see a gap in the severity of concept resurgence between the fine-tuning algorithms and those using closed-form edits. Specifically, UCE is quite robust, exhibiting very small resurgence across tasks. We conjecture that UCE is the strongest type of closed-form edit, as it modifies the cross attention weights to directly map the target concept to a higher-level (more abstract) concept. For example, if the target concept is a particular celebrity, it may be mapped to the more abstract concept like “a Person" or “a Celebrity". In contrast, MACE modifies the cross-attention weights to map the embeddings of all the surrounding words in the given prompts to be similar to embeddings of the surrounding words after replacing the target concept with a more abstract one. This difference means that MACE does not directly optimize the parameter update to move the target concept embedding towards the abstract concept embedding. Furthermore, because MACE incorporates unlearning the target concept information via fine-tuning, this might leave it more vulnerable to concept resurgence than UCE, which is based on a direct closed-form edit.
Parameter Choice The second algorithmic factor we examine is which subsets of parameters are updated in the unlearning phase, and which (potentially overlapping) subsets of parameters are further fine-tuned. We start by showing how these choices potentially explain why UCE is more robust to concept resurgence than the other three algorithms. As discussed above, UCE only modifies the cross-attention weights with a closed form edit. As discussed in [14], this approach is very effective for concepts that are localized to the words themselves (e.g. the name of a celebrity; contrast this to unsafe content, which is a more abstract concept). Applying LoRA fine-tuning after UCE unlearning, we find no evidence of concept resurgence. We then instead fine-tune the full set of parameters, which yields a small degree of resurgence. Finally, motivated by this result, we choose to fully fine-tune the cross-attention layers only. We see that the resurgence is comparable between the two (Table 4), suggesting that the nature of UCE’s closed-form edit being localized to the cross-attention layers may make it very robust.
The second difference between the four algorithms is the subset of model parameters that are updated in the unlearning step. Section 3 focuses primarily on modifying the cross-attention layers (with the exception of MACE, which also updates the rest of the model parameters via LoRA fine-tuning). Here, we focus on ESD in the single celebrity erasure task and the copyright erasure task, which both exhibit very high degrees of concept resurgence. In each of these tasks, we vary vary the subset of parameters that are updated in the unlearning step: either all of the parameters, all of the parameters except those in the cross-attention layers, and only those in the cross-attention layers. We find that the cross-attention parameters do indeed play the most important role in unlearning for these tasks and that unlearning on all the parameters only provided marginal gains in preventing resurgence (Fig. 4).
5 Why Does Concept Resurgence Occur?
Finally, to better understand the root cause(s) of concept resurgence, we explore this phenomenon in a simplified one-dimensional setting, where the distribution of interest is a simple mixture of standard Gaussians. This (intentionally stylized) model provides valuable intuition and insight into the dynamics of concept resurgence.
Setup. We construct two Gaussian distributions and with means and respectively. We fix for both distributions. will model the original “pretraining" distribution, and will model the distribution on which the model is fine-tuned. Next, we define a “concept" as the following membership function (i.e. a concept is represented as an interval on the real line). We model exact unlearning as (re)training a diffusion model on the data sampled from the original distribution, excluding values from the interval . To construct this distribution, we simply perform rejection sampling from , rejecting any samples which fall in the interval .
In this setting, we model an “approximate unlearning" algorithm as one which approximates the desired data distribution but leaves a probability mass of on the unwanted interval. indicates exact unlearning and indicates no unlearning. We model approximate unlearning by simply training on a sample of data from where we performed modified rejection sampling with a tolerance parameter of — if a sample lies in , it is rejected with probability ; otherwise, it is retained with probability .
This setup allows us to investigate how varying levels of probability mass which remain in the unlearned concept region — corresponding to varying degrees of “success" in the initial unlearning step — can lead to concept resurgence. For the sake of this example, we start with a baseline level of approximate unlearning quality at . Finally, after applying unlearning, we fine-tune the resulting model on data sampled from (after first rejecting any samples which lie in ).
Training. With this setup, we train denoising score matching models (the same techniques used in the Stable Diffusion models studied previously) to model these distributions. Our diffusion models are based on the variance exploding SDE, where we choose the diffusion coefficient to be . We train separate diffusion models for each value of to represent varying unlearning quality. Afterwards, we fine-tune each of these models on samples from . We use a KL divergence penalty in the score denoising loss when fine-tuning to prevent catastrophic forgetting.
The experiments we present are for the following setup: , and .
Evaluation. To measure resurgence in this setting, we measure the average log-likelihood of five equally spaced points in the unwanted concept interval and the number of samples generated by the diffusion model that contain the concept. In practice, it is intractable to compute the log-likelihood of data under the learned distribution. However, because our model is one-dimensional, we can approximately compute the log-likelihoods via numerical integration. We provide more details about this procedure in Appendix G.
Results. We first consider exact unlearning, i.e. . We plot the distributions of the original samples from both and with the learned distributions after (1) exact unlearning and (2) fine-tuning after exact unlearning in Figure 5. It is important to note that even under exact unlearning, the diffusion model leaves some non-zero probability mass on the unlearned concept region. We conjecture this is due to the implicit bias of diffusion models for learning smooth distributions (as also characterized by other works [1]), which leads to some mass being placed on the unwanted concept interval even though this region is outside the support of the training data. We further observe that fine-tuning amplifies this small amount of additional probability mass on the unlearned concept interval.
We now examine how this phenomenon changes as a function of the amount of probability mass which remains in the unwanted concept interval after unlearning. The degree of resurgence (as measured by the number of samples that contain the concept) is constant as we increase this probability (Figure 6). The average log likelihood also increases after fine-tuning, suggesting that one cause for resurgence might be the model’s inductive bias towards learning smooth distributions, which in turn places some small probability mass on the unwanted concept interval.
Although this mass may be negligible — so small that it is difficult to detect with sampling-based approaches, subsequent fine-tuning can lead to significant concept resurgence. This model is consistent with our empirical results, as well as those which appear elsewhere in the literature [13, 22, 13, 27] — unlearning algorithms typically suppress (rather than fully remove) the probability of generated an unwanted concept. Of course, this stylized model does not capture the full complexity of modern text-to-image diffusion models like Stable Diffusion v1.4 and v2.1, but our results shed light on possible factors driving concept resurgence and suggest avenues for future work.
@@ 6 Discussion and Limitations
The scale of generative models introduces new challenges, including the risk of learning concepts that are unsuitable or undesirable for certain downstream applications.
Ideally, unlearning algorithms would enable the precise and permanent removal of unwanted concepts while preserving the model’s overall utility. Reality, however, is more complex.
@@
Our investigation opens up several important avenues for future work. For example, we do not provide a theoretical characterization of concept resurgence, nor do we present a strategy designed to prevent it from happening. Both developments could help to enhance the robustness of unlearning methods. Additionally, our evaluations focus on a mix of well-known academic benchmarks and synthetic tasks, and further research is necessary to assess the prevalence of concept resurgence in real-world deployments (particularly the effect of interleaving a large number of compositional updates, which may exacerbate these vulnerabilities).
@@
Appendix A Additional Qualitative Examples
In this section we include qualitative results for the copyright and object erasure tasks in Figure 7 and Figure 8, respectively. These results are analogous to those presented in Figure 2 for the celebrity erasure task. We choose to exclude qualitative examples of resurgence for the unsafe content task, as these may be upsetting. For a quantitative evaluation of this task across all four unlearning algorithms, we refer readers to Table 1.
Appendix B Unlearning Tasks
For the copyright task, we choose the concept “Iron Man" to erase. We generate five prompts that we provide the model with five different random seeds to evaluate its knowledge of Iron Man. These prompts were:
1.
“A photo of [name]"
2.
“A portrait of [name]"
3.
“A pose of [name] in action"
4.
“An illustration of [name]"
5.
“An iconic scene of [name]"
Additionally, we create retain set of copyright characters that include: Albus Dumbledore, Anna, Aquaman, Aragorn, Arwen, Barbie, Bart Simpson, Batman, Black Panther, Black Widow, Boromir, Bugs Bunny, Buzz Lightyear, C-3PO, Captain America, Catwoman, Chewbacca, Daffy Duck, Darth Vader, Doctor Strange, Donald Duck, Darth Vader, Doctor Strange, Donald Duck, Donkey, Dr. Watson, Draco Malfoy, Dracula, Ebenezer Scrooge, Elsa Mars, Elsa, Fiona, Flash, Frankenstein’s Monster, Fred Flinstone, Frodo Baggins, Galadriel, Gandalf, Gollum, Goofy, Green Lantern, Hagrid, Han Solo, Harley Quinn, Harry Potter, Hermione Granger, Homer Simpson, Huckleberry Finn, Hulk, Jack Sparrow, Joker, Juliet, Katniss Everdeen, Kirby, Kylo Ren, Lara Croft, Legolas, Lex Luthor, Link, Loki, Luigi, Luke Skywalker, Luna Lovegood, Mario, Master Chief, Mickey Mouse, Minnie Mouse, Moana, Neo, Neville Longbottom, Obi-Wan Kenobi, Oliver Twist, Patrick Star, Peter Griffin, Pikachu, Princess Leia, Princess Peach, R2D2, Romeo, Ron Weasley, Samwise Gamgee, Sauron, Scarlet Witch, Scooby-Doo, Severus Snape, Shaggy, Sherlock Holmes, Shrek, Simba, Snoopy, Sonic the Hedgehog, Spider-Man, Spongebob Squarepants, Superman, Thor, Tom Sawyer, Tony Montana, Voldemort, Willy Wonka, Wonder Woman, Woody, and Yoda.
For the unsafe content task, we select a subset of concepts from the Inappropriate Images Prompts (I2P) [37] dataset. We are focused on erasing the concepts hate, self-harm, violence, and harassment. We select prompts labeled as such in the dataset and that have a score of at least 70% or more on the Q16 percentage. This percentage represents how many times out of 10 samples the Q16 classifier classified the image as inappropriate.